OpenStudy (anonymous):

On Wall Street, a stock that shoots up in price is said to grow parabolically. An internet stock called Fibonacci.com has a price given by f (x) = 3x 2 − 5x + 6, where x is the number of the days the company has been in business. What is the price of Fibonacci.com on the 4th day?

OpenStudy (anonymous):

f(4) = 3(4)2 - 5 (4) + 6 = 3*16 - 20 + 6 = 48 - 14 = 34.